# Fostering Collective Discourse: A Distributed Role-Based Approach to Online News Commenting

### v1

@@ 3.4.2. Summarizing
After clusters are created per the process described above, LV1 users can summarize accepted clusters. Upon clicking the “Summarize” button in the cluster, a popup modal appears with an AI-suggested summary (Figure 4). While users can use the suggested summary, they are encouraged to revise or create their own to critically engage with the discussion, ensuring that the final summary reflects a comprehensive range of viewpoints. Once a LV1 user has created a summary, it will appear at the top of the cluster and be visible to all users. After a cluster has been summarized, no additional comments can be added, as this would affect the appropriateness of the existing summary. Thus, when creating summaries, LV1 users are instructed to check that all relevant perspectives are included in the cluster, and assess if there is room for new comments. The summary of the cluster is displayed under the discussion thread on the first page, thereby informing users of the key discussions emerging in each thread.
Figure 4. Workflow of Summarizing: (1)LV1 users summarize accepted clusters by clicking the ’Summarize’ button within the cluster. (2) A modal will display an AI-generated summary suggestion, which users can revise or replace with their own (left). (3) Once finalized, the summary becomes visible to all users and is displayed at the top of the cluster (right).:
@@ 3.4.3. Threading
New discussion thread topics are created through the suggestions and review activities of LV2 users, as outlined in Figure 5. As discussions within each topic evolve, LV2 users can propose new topics for threads by clicking the “Suggest New Thread” button. A popup modal will appear with an AI-generated topic suggestion and a guiding question for the new discussion, designed to assist users in formulating new topics. These topics and guiding questions are generated based on the article’s content. Users have the option to select the AI-suggested topic by checking the box, but they are encouraged to suggest a topic that incorporates and aligns with the ongoing discussion flow.
To review these created threads, users click the “Review Threads” button. A popup modal will display a list of suggested topics from other users, which can be approved or declined via checkboxes. A new discussion thread is created once a topic is approved by the required number of LV2 users. We set this requirement to three participants for the purpose of the user study, but it can be adjusted to different thresholds. The newly accepted threads will appear at the bottom of the original thread boxes on the first page. All users can access these newly created threads and participate in the discussion.
Figure 5. Workflow of Threading: (1) LV2 users propose new thread topics by selecting from AI-suggested topics or by suggesting their own. (2) Users review these topics by approving or declining each proposed thread. (3) A new discussion thread is created and becomes visible to all users once a topic is approved by the required number of LV2 users.:
@@ 4.1. Participants
We recruited a total of 40 participants to use the commenting system browser extension through communities managed by our institution. We asked them to fill out a survey asking about the frequency of reading news articles and writing comments and their motivations to do those activities. We filtered out the participants who never read articles during the week to verify our system among those who regularly consume news articles. Two participants dropped out due to scheduling constraints, so a total of 38 participants (Age=19-31, $M=23.12$, $SD=3.25$; 22 male and 16 female) were included in our study. Participants received 40, [^0] KRW for their participation over six days, with an expected usage of 10 to 20 minutes of the system each day.
@@ 4.2.1. System Usage
Each participant was assigned a specific level prior to the start of the user study, and this level remained constant throughout the experiment. We conducted a within-subject study comparing our system with a baseline. For the baseline, we implemented a system that displays the standard commenting section, featuring only the commenting, reply, edit, delete and like functions that users are familiar with. The order in which participants used the baseline system and our system was counterbalanced.
@@
Participants were divided into a total of six groups, balancing the number of participants at each level. Each group started at a different time, with a 12-hour interval between the start time. Since each level of participants has different needs to facilitate discussion in the system, the goal of assigning groups with different participation time periods was to have all levels work simultaneously. The number of participants at each level was adjusted based on the participation levels and ongoing activities observed during the pilot study. Figure 6 summarizes the study procedure including group assignments, level distribution, article sequence and staggered participation schedule. After using the system, we asked participants to complete a post-survey about their general experiences and the impact of using our system. The post-survey questions are presented in Appendix B.
Figure 6. Procedure of the study. We conducted a within-subject study over six days for each participant, using both our system and a traditional commenting system, with the order balanced. The study included six articles, two for each of three different topics, with the same topics spaced evenly. Participants were divided into six groups with staggered start times to have all levels function simultaneously. The number of levels assigned to each group is shown on the right.:
@@ 4.3. Measures and Hypotheses
We derived the following hypotheses by linking these measures to the design intentions of our system. The structured environment was expected to encourage more active but concise participation (H1), while clustering and summarization features were designed to surface a broader distribution of perspectives (H2). Guided prompts and distributed roles were intended to strengthen reasoning processes (H3), whereas structured participation was anticipated to temper negative affect while maintaining prosocial tone (H4, H5). Together, these hypotheses capture our expectation that the system would foster more balanced, civil, and analytically grounded discussion.
@@
Comments in the system condition will demonstrate higher argument quality, with a greater proportion of claims supported by evidence.
@@ 5.1.1. User Engagement Metrics
Since participants contributed comments in both conditions, we analyzed likes and comment length using multilevel models to account for the non-independence of observations(Table 2). Specifically, we fit generalized linear mixed models (NB-GLMMs) with condition (baseline and system) as a fixed effect and random intercepts for participants. We adopted this model since the diagnostic checks indicated overdispersion, with the variance of the data exceeding the mean. Comments in the system condition received fewer likes on average (IRR = 0.82, 95% HDI [0.56, 1.15]), but this difference was not significant. For word count, system comments were significantly shorter (IRR = 0.74, 95% HDI [0.66, 0.82]), corresponding to an approximate 26% reduction relative to baseline. These findings suggest that while the system encouraged more frequent and concise contributions, it did not significantly alter patterns of peer endorsement.
@@
Table 2. Results of negative binomial generalized linear mixed models (NB-GLMMs) comparing system and baseline conditions. IRRs (Incidence Rate Ratios) less than 1 indicate reductions in the system condition relative to baseline. System comments were significantly shorter than baseline comments, while differences in like counts were not significant.:
@@ 5.1.2. Details of System Usage
Clustering Each article demonstrated a similar trend, with 5 to 7 finalized clusters emerging as the outcome of extensive clustering activities. Notably, the formation of these completed clusters required 16 to 36 total individual clustering activities per article, highlighting the intensive review and refinement process involved. Pending clusters, which had not yet reached final approval, predominantly exhibited an Accept(2/3) status, suggesting that many were close to completion. For example, in the “Tech” article, 10 of the 12 pending clusters were nearing acceptance, while the “Crime” article showed 8 out of 12 pending clusters in a same state. The “Economy” article displayed a smaller scale of activity, with both of its 2 pending clusters marked as Accept(2/3). The average of 9.67 activities resulted in accepted clusters, while 9.33 activities concluded with denial, indicating significant deliberation and evaluation among LV1 users during the clustering process. These findings highlight that, although the finalized clusters represent a smaller subset of the total activities, the dynamics between suggestions from LV0 contributors and evaluations by LV1 reviewers created a balanced and constructive tension.
@@
[…]
Table 3. Summary of system usage statistics, including the number of clusters, summaries, and threads created by users. A breakdown of activities for creating clusters and threads is provided, detailing the number of review activities conducted for these creations. The total number of activities is presented, divided into accepted, pending, and denied for both clusters and threads.:
@@
Table 4. Thread topics for the three articles, shown left to right as: (1) initial topics provided by the system, (2) user-generated topics accepted by other users, and (3) pending topics under review. AI-suggested topics are indicated in italics.:
@@
Table 5. Summary of key outcome measures across conditions. Baseline mean reflects the intercept; $\Delta$ indicates the change under the system condition. Significant effects are bolded.:
@@ 5.2.1. Perspective Diversity
To assess whether the system fostered greater diversity of perspectives within discussions, we embedded all comments using all-MiniLM-L6-v2 (Sentence-BERT) and clustered them into semantic groups using K-means. For each article–condition pair, we computed two standard entropy-based measures: normalized entropy (H_norm), capturing the evenness of distribution across clusters, and coverage, reflecting the proportion of clusters represented. Mixed-effects linear models were fit with condition as a fixed effect and article as a random intercept. Results showed that H_norm was significantly higher in the system condition ($\beta$ = +0.070, 95% CI [0.012, 0.129], z = 2.38, p = .018), indicating more balanced participation across perspectives. By contrast, coverage did not significantly differ between conditions ($\beta$ = +0.033, 95% CI [-0.032, 0.099], z = 1.00, p = .317). These findings suggest that the system primarily enhanced the evenness of perspectives expressed, reducing dominance of a few clusters, but did not expand the overall range of distinct viewpoints.
@@ 5.2.2. Argument Strength
We next examined argument strength using the Supported-Claim Ratio (SCR), defined as the proportion of claims accompanied by at least one supporting premise or piece of evidence. Sentences were labeled (claim, premise, evidence, other) via zero-shot classification with facebook/bart-large-mnli, and SCR was computed at the level of individual participants within conditions. Mixed-effects linear models with participant random intercepts and article variance components indicated a baseline mean of 0.228 and 0.153 under system, yielding a significant decrease ($\beta$ = -0.075, SE = 0.027, 95% CI [-0.128, -0.021]). These findings suggest that although the system may have stimulated participants to produce more claims overall, these claims were less frequently supported with explicit reasoning or evidence. This pattern points to a trade-off: system support broadened engagement in claim-making but at the expense of argumentative robustness.
@@ 5.2.3. Emotion
To examine the effect of the system on emotional expression, we classified sentence-level probabilities for 29 fine-grained emotions. From these outputs, we derived comment-level measures: (a) emotionality (1 - P(neutral)) and (b) probabilities for six core emotions (anger, joy, sadness, fear, disgust, surprise). Mixed-effects logistic models with participant random intercepts were fit, with outcomes logit-transformed. Results indicated a significant reduction in overall emotionality under system support (baseline = 0.656, system = 0.552, $\beta$ = -0.109, SE = 0.019, 95% CI [-0.147, -0.071], p ¡ .001). Core emotion analyses revealed significant decreases in joy ($\beta$ = -0.529, p ¡ .001), sadness ($\beta$ = -0.590, p ¡ .001), fear ($\beta$ = -0.512, p ¡ .001), and surprise ($\beta$ = -0.383, p ¡ .01), while anger and disgust showed small, non-significant increases. Taken together, these findings suggest that system assistance dampened emotional expression across comments, yielding a more neutral and analytical tone. Both positive and negative emotions were suppressed, while antagonistic emotions such as anger and disgust remained unchanged, indicating that the system primarily reduced emotionality without amplifying hostility.

### v2

@@ 3.4.2. Summarizing
After clusters are created per the process described above, LV1 users can summarize accepted clusters. Upon clicking the “Summarize” button in the cluster, a pop-up modal appears with an AI-suggested summary (Figure 4). While users can use the suggested summary, they are encouraged to revise or create their own to critically engage with the discussion, ensuring that the final summary reflects a comprehensive range of viewpoints. Once an LV1 user has created a summary, it will appear at the top of the cluster and be visible to all users. After a cluster has been summarized, no additional comments can be added, as this would affect the appropriateness of the existing summary. Thus, when creating summaries, LV1 users are instructed to check that all relevant perspectives are included in the cluster, and assess if there is room for new comments. The summary of the cluster is displayed under the discussion thread on the first page, thereby informing users of the key discussions emerging in each thread.
Figure 4. Workflow of Summarizing: (1) LV1 users summarize accepted clusters by clicking the ‘Summarize’ button within the cluster. (2) A modal will display an AI-generated summary suggestion, which users can revise or replace with their own (left). (3) Once finalized, the summary becomes visible to all users and is displayed at the top of the cluster (right).
The figure shows the workflow for summarizing accepted clusters in a system by Level 1 (LV1) users. On the left, there is a ‘Summarize’ button within a cluster for LV1 users. After clicking it, a modal appears displaying an AI-generated summary suggestion. The suggestion says: ’The comments suggest media highlights drug risks through celebrity cases, but law enforcement must treat all equally to deter drug issues.’ Users can revise or replace the AI-generated summary with their own text by typing in the provided field. On the right side, the final summary is displayed at the top of the clustered comments once it is saved. Below the summary, users’ original comments that were included in the clusters are visible. The figure includes step-by-step instructions for this process: (1) LV1 users click the ’Summarize’ button, (2) a modal with the AI-suggested summary appears, allowing the user to edit or replace it with their own summarization, and (3) the finalized summary becomes visible to all users and is displayed at the top of the cluster.
@@ 3.4.3. Threading
New discussion thread topics are created through the suggestions and review activities of LV2 users, as outlined in Figure 5. As discussions within each topic evolve, LV2 users can propose new topics for threads by clicking the “Suggest New Thread” button. A pop-up modal will appear with an AI-generated topic suggestion and a guiding question for the new discussion, designed to assist users in formulating new topics. These topics and guiding questions are generated based on the article’s content. Users have the option to select the AI-suggested topic by checking the box, but they are encouraged to suggest a topic that incorporates and aligns with the ongoing discussion flow.
To review these created threads, users click the “Review Threads” button. A pop-up modal will display a list of suggested topics from other users, which can be approved or declined via checkboxes. A new discussion thread is created once a topic is approved by the required number of LV2 users. We set this requirement to three participants for the purpose of the user study, but it can be adjusted to different thresholds. The newly accepted threads will appear at the bottom of the original thread boxes on the first page. All users can access these newly created threads and participate in the discussion.
Figure 5. Workflow of Threading: (1) LV2 users propose new thread topics by selecting from AI-suggested topics or by suggesting their own. (2) Users review these topics by approving or declining each proposed thread. (3) A new discussion thread is created and becomes visible to all users once a topic is approved by the required number of LV2 users.
The figure shows the workflow for threading in a system where Level 2 (LV2) users propose and review new discussion threads. On the left side, LV2 users are presented with the option to ’Suggest New Thread’ by either selecting an AI-suggested topic or writing their own. The AI-suggested topic shown is ’Responsibility of Tech Companies’, with the supporting question: ’Should tech companies be held accountable for the misuse of AI tools like Grok on their platforms?’ Users can write their own new topic and add a supporting question in the provided fields. In the middle of the figure, a ’Review Suggested Threads’ section allows LV2 users to approve or decline pending topics proposed by others. Two topics are shown: ’Responsibility of Tech Companies’ and ’Positive Applications of Grok’, with checkmark and cross icons for approval or rejection. On the right side, once a thread is accepted by the required number of LV2 users (N=3), the new discussion thread appears under the ’Discussions About the Article’ section. The newly created thread, titled ’Responsibility of Tech Companies’, is now visible to all users. The figure includes step-by-step instructions for this process: (1) LV2 users suggest new thread topics, (2) review suggested topics by other LV2 users, and (3) a new thread is created and becomes visible once it is approved by the required number of LV2 users.
@@ 4.1. Participants
We recruited a total of 40 participants to use the commenting system browser extension through communities managed by our institution. We asked them to fill out a survey asking about the frequency of reading news articles and writing comments and their motivations to do those activities. We filtered out the participants who never read articles during the week to verify our system among those who regularly consume news articles. All study tasks, including reading news articles and writing comments or summaries, were conducted in English. For recruitment, we specified that participation required the ability to understand English news content and produce written responses in English, and participants confirmed that they were able to do so as part of the screening survey. Participants used English for all written interactions within the system. Two participants dropped out due to scheduling constraints, so a total of 38 participants (Age=19-31, $M=23.12$, $SD=3.25$; 22 male and 16 female) were included in our study. Participants received 40, [^0] KRW for their participation over six days, with an expected usage of 10 to 20 minutes of the system each day.
@@ 4.2.1. System Usage
Each participant was assigned a specific level prior to the start of the user study, and this level remained constant throughout the experiment. We conducted a within-subject study comparing our system with a baseline. For the baseline, we implemented a system that displays the standard commenting section, featuring only the commenting, reply, edit, delete, and like functions that users are familiar with. The order in which participants used the baseline system and our system was counterbalanced.
@@
Participants were divided into a total of six groups, balancing the number of participants at each level. Each group started at a different time, with a 12-hour interval between the start times. Since each level of participants has different needs to facilitate discussion in the system, the goal of assigning groups with different participation time periods was to have all levels work simultaneously. The number of participants at each level was adjusted based on the participation levels and ongoing activities observed during the pilot study. Figure 6 summarizes the study procedure including group assignments, level distribution, article sequence and staggered participation schedule. After using the system, we asked participants to complete a post-survey about their general experiences and the impact of using our system. The post-survey questions are presented in Appendix B.
Figure 6. Procedure of the study. We conducted a within-subject study over six days for each participant, using both our system and a traditional commenting system, with the order balanced. The study included six articles, two for each of three different topics, with the same topics spaced evenly. Participants were divided into six groups with staggered start times to have all levels function simultaneously. The number of levels assigned to each group is shown on the right.
The figure displays the procedure of a within-subject study conducted over seven days, involving six participant groups. Each group alternates between using the system (’System’ in blue) and a traditional commenting system (’Baseline’ in yellow) across six days. The columns represent days, and each row represents a group. Groups 1 to 3 begin with the system, while groups 4 to 6 begin with the traditional commenting system, with alternating use on different days. The system and baseline are evenly distributed across days to ensure balance. On the right side of the figure, the number of users assigned to each level (LV0, LV1, LV2) is shown for each group. Group 1 include 2 LV0, 2 LV1 users, group 2 include 2 LV0, 4 LV1, 2 LV2 users, and group 3 include 2 LV0, 3 LV1, 2 LV2 users. Similarly, groups 4 to 6 also show the same distribution as groups 1 to 3, with a total of 38 participants participating in the study. At the bottom, a color-coded key shows the three topics discussed during the study: A (Tech), B (Crime), and C (Economy), with these topics spaced evenly throughout the study. The six groups have staggered start times to make all levels function simultaneously.
@@ 4.3. Measures and Hypotheses
We derived the following hypotheses by linking these measures to the design intentions of our system. First, our system reduces the need for participants to individually reconstruct the shared context of the discussion. By presenting a pre-structured environment, where related ideas are already grouped and prior contributions are visible in an organized manner, the system eliminates redundant cognitive work typically required in free-form commenting. Guided prompts further streamline expression by helping participants focus on contributing new information or clarifying ideas. As a result, we expect that users can participate more frequently yet more concisely, without compromising the clarity or perceived value of their comments.
@@
Second, clustering and summarization improve the overall visibility of distinct viewpoints. By organizing contributions into a clear structure, the system makes the range of perspectives easier to perceive and navigate, which we expect will support broader engagement with diverse ideas.
@@
Third, guided prompts and distributed roles support users in constructing clearer and more analytically grounded arguments. By directing attention toward justification and evidence, the system encourages reasoning that is more explicit and well-supported, therefore improving argument strength.
@@
Comments in the system condition will demonstrate higher argument strength, with a greater proportion of claims supported by evidence.
Fourth, the system guides participation toward task-oriented interactions, which are expected to reduce opportunities for unstructured, reactive exchanges that often escalate negativity. By providing clear contribution pathways and shared conversational targets, users can focus on collaborative meaning-making rather than interpersonal conflict. We therefore expect emotional expression to remain constructive, with increased politeness in user interactions.
@@
Together, these hypotheses capture our expectation that the system would foster more balanced, civil, and analytically grounded discussion.
@@ 5.1.1. User Engagement Metrics
Since participants contributed comments in both conditions, we analyzed likes and comment length using multilevel models to account for the non-independence of observations(Table 2). Specifically, we fit generalized linear mixed models (NB-GLMMs) with condition (baseline and system) as a fixed effect and random intercepts for participants. We adopted this model since the diagnostic checks indicated overdispersion, with the variance of the data exceeding the mean. Comments in the system condition received fewer likes on average (IRR = 0.82, 95% HDI [0.56, 1.15]), but this difference was not significant. For word count, system comments were significantly shorter (IRR = 0.74, 95% HDI [0.66, 0.82]), corresponding to an approximate 26% reduction relative to baseline. These findings suggest that while the system encouraged more frequent and concise contributions, it did not significantly alter patterns of peer endorsement. This result is consistent with H1, indicating that the structured environment increased active but concise participation without reducing the perceived value of comments.
@@
This table compares engagement metrics between the system and baseline across three article topics. It shows an increase in commenting activity for all topics, with 61, 67, and 61 comments for the baseline, and 89, 79, and 62 comments for the system (in order of tech, crime, and economy). The levels of replies and likes remained similar across all topics. Additionally, the average comment length and word count were significantly shorter for all three topics, with an average word count of 49.13 for the baseline and 34.96 for the system.
@@
This table compares comments made in the system condition with those in the baseline condition. It reports two measures: the number of likes and the length of comments. On average, comments in the system condition received about 18% fewer likes than baseline comments, but this difference was not statistically meaningful. In contrast, system comments were about 26% shorter than baseline comments, and this reduction was statistically meaningful.
Table 2. Results of negative binomial generalized linear mixed models (NB-GLMMs) comparing system and baseline conditions. IRRs (Incidence Rate Ratios) less than 1 indicate reductions in the system condition relative to baseline. System comments were significantly shorter than baseline comments, while differences in like counts were not significant.
@@ 5.1.2. Details of System Usage
Clustering Each article demonstrated a similar trend, with 5 to 7 finalized clusters emerging as the outcome of extensive clustering activities. Notably, the formation of these completed clusters required 16 to 36 total individual clustering activities per article, highlighting the intensive review and refinement process involved. Pending clusters, which had not yet reached final approval, predominantly exhibited an Accept(2/3) status, suggesting that many were close to completion. For example, in the “Tech” article, 10 of the 12 pending clusters were nearing acceptance, while the “Crime” article showed 8 out of 12 pending clusters in the same state. The “Economy” article displayed a smaller scale of activity, with both of its 2 pending clusters marked as Accept(2/3). The average of 9.67 activities resulted in accepted clusters, while 9.33 activities concluded with denial, indicating significant deliberation and evaluation among LV1 users during the clustering process. These findings highlight that, although the finalized clusters represent a smaller subset of the total activities, the dynamics between suggestions from LV0 contributors and evaluations by LV1 reviewers created a balanced and constructive tension.
@@
Tech 6 36 11 12 13 6 4 3 1 0
Crime 7 31 11 12 8 2 4 3 1 0
Economy 5 16 7 2 7 5 3 1 2 0
Table 3. Summary of system usage statistics, including the number of clusters, summaries, and threads created by users. A breakdown of activities for creating clusters and threads is provided, detailing the number of review activities conducted for these creations. The total number of activities is presented, divided into accepted, pending, and denied for both clusters and threads.
This table presents a summary of our system usage statistics, separated by clusters, summaries, and threads. It shows the number of clusters, summaries, and threads created by users across all three articles. For clustering and threading activities, a breakdown is provided, detailing the number of review activities conducted for these creations. The total number of activities is divided into accepted, pending, and denied for both clusters and threads. On average, 6 clusters, 4.3 summaries, and 2.3 threads were created per article. For clustering activities, an average of 27.7 clustering activities were conducted to create the total number of clusters, having similar levels of accepted and denied clustering activities.
@@
Table 4. Thread topics for the three articles, shown left to right as: (1) initial topics provided by the system, (2) user-generated topics accepted by other users, and (3) pending topics under review. AI-suggested topics are indicated in italics.
This table lists thread topics for Technology, Crime, and Economy articles. For each article, columns present the system-provided initial topics, user-created topics that were accepted, and topics still pending; AI-suggested items are marked in italics.
@@
Table 5. Summary of key outcome measures across conditions. Baseline mean reflects the intercept; $\Delta$ indicates the change under the system condition. Significant effects are bolded.
The table shows the summary of the key outcome measures across the conditions, along the four measures of perspective diversity, argument strength, emotion, and politeness. The baseline mean reflects the intercept and the change under the system condition is denoted as delta. We observed significant results in all measures except for emotions of anger and disgust, and politeness.
@@ 5.2.1. Perspective Diversity
To assess whether the system fostered greater diversity of perspectives within discussions, we embedded all comments using all-MiniLM-L6-v2 (Sentence-BERT) and clustered them into semantic groups using K-means. For each article–condition pair, we computed two standard entropy-based measures: normalized entropy (H_norm), capturing the evenness of distribution across clusters, and coverage, reflecting the proportion of clusters represented. Mixed-effects linear models were fit with condition as a fixed effect and article as a random intercept. Results showed that H_norm was significantly higher in the system condition ($\beta$ = +0.070, 95% CI [0.012, 0.129], z = 2.38, p = .018), indicating more balanced participation across perspectives. By contrast, coverage did not significantly differ between conditions ($\beta$ = +0.033, 95% CI [-0.032, 0.099], z = 1.00, p = .317). Taken together, these results provide partial support for H2: the system improved the evenness of perspectives expressed by reducing the dominance of a few perspectives, though it did not increase the total number of distinct viewpoints expressed.
@@ 5.2.2. Argument Strength
We next examined argument strength using the Supported-Claim Ratio (SCR), defined as the proportion of claims accompanied by at least one supporting premise or piece of evidence. Sentences were labeled (claim, premise, evidence, other) via zero-shot classification with facebook/bart-large-mnli, and SCR was computed at the level of individual participants within conditions. Mixed-effects linear models with participant random intercepts and article variance components indicated a baseline mean of 0.228 and 0.153 under system, yielding a significant decrease ($\beta$ = -0.075, SE = 0.027, 95% CI [-0.128, -0.021]). Accordingly, these results do not support H3, indicating that the structured participation did not translate into stronger explicit evidence-based reasoning.
@@ 5.2.3. Emotion
To examine the effect of the system on emotional expression, we classified sentence-level probabilities for 29 fine-grained emotions. From these outputs, we derived comment-level measures: (a) emotionality (1 - P(neutral)) and (b) probabilities for six core emotions (anger, joy, sadness, fear, disgust, surprise). Mixed-effects logistic models with participant random intercepts were fit, with outcomes logit-transformed. Results indicated a significant reduction in overall emotionality under system support (baseline = 0.656, system = 0.552, $\beta$ = -0.109, SE = 0.019, 95% CI [-0.147, -0.071], p ¡ .001). Core emotion analyses revealed significant decreases in joy ($\beta$ = -0.529, p ¡ .001), sadness ($\beta$ = -0.590, p ¡ .001), fear ($\beta$ = -0.512, p ¡ .001), and surprise ($\beta$ = -0.383, p ¡ .01), while anger and disgust showed small, non-significant increases. Taken together, these findings suggest that system assistance dampened emotional expression across comments, yielding a more neutral and analytical tone. Both positive and negative emotions were suppressed, while antagonistic emotions such as anger and disgust remained unchanged, indicating that the system primarily reduced emotionality without amplifying hostility. Thus, these results offer partial support for H4: the system discouraged negative affect without promoting positive expression.